OpenStudy (anonymous):

six sides of cube are marked as follows: three sides with number -3; two sides with number 2; and one side with number 0. a) what is the probability that in 4 tries maximum once you get number -3 (i solved this one and got 1/16) b) what is the probability that in 3 throws you get that sum of numbers is negative c) how many times you must throw dice so that you get 0 once with 0,999 probability d) cube is repeatedly thrown with rules: - if you get negative number nothing happens - if you get 0 or 2, you replace that number with -3 you throw so long until all sides of cube

OpenStudy (anonymous):

d) ...you throw so long until all sides of cube are -3. what is the probability that it would take more than 3 throws to achieve that. this is for my friend, its been long time i did tasks like this, so i hoped you could help... tnx

OpenStudy (anonymous):

a- 15/16. probability of that you don't get number -3 is 1/2. if you don't get number -3 four times, (1/2)*(1/2)*(1/2)*(1/2)=1/16. 1-(1/16)=15/16

OpenStudy (anonymous):

yes but probability that you only once get -3... probability to get -3 = 1/2 probability not to get -3 = 1/2 so: (1/2)^4=1/16